# A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability

### v2

@@ 4.3. Privacy Features
Privacy in digital payment systems is often defined through different sub-properties focusing on anonymity of participants, value confidentiality, plausible deniability, and unlinkability (e.g., (Almashaqbeh and Solomon, 2022; Amarasinghe et al., 2019; Bernal Bernabe et al., 2019; Genkin et al., 2018)).
@@ 4.3.1. Sender Anonymity
Asymmetric anonymity indicates that either the sender’s or receiver’s identity is protected (e.g., (Chaum, 1983; Sander and Ta-Shma, 1999; Miers et al., 2013; Dold, 2019; Zhang et al., 2019; BIS Innovation Hub, 2023; Goodell et al., 2023; Yu et al., 2024; Androulaki and Karame, 2014)). Sender anonymity focuses only on hiding the payer’s identity, usually using the following techniques: blind signatures (e.g., (Chaum, 1983; Dold, 2019; Goodell et al., 2023; Liu et al., 2022)) and their variants—one-time (e.g., (Hayes, 1990)), threshold (e.g., (Sarencheh et al., 2025; Kiayias et al., 2022; Androulaki et al., 2020)), and randomizable (e.g., (Tomescu et al., 2022; Wüst et al., 2019))—, ring signatures (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), group signatures (e.g., (Zhang et al., 2019)), and commitments with ZKPs (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)).
@@
Chaum (Chaum, 1983) proposes blind signature to unlink the sender request of token issuance and its later spending. To reveal the sender on double spending, Hayes (Hayes, 1990) introduces one-time blind signatures. However, since each transaction requires a new signing key, the storage and computation management is cumbersome; moreover, one-time signatures often require large signature size and expensive verification, making them impractical for high-throughput systems. UTT (Tomescu et al., 2022) implements blind signatures using randomizable signatures (Pointcheval and Sanders, 2016), which allow signatures to be altered using fresh randomness without compromising their validity. All these schemes rely on a central issuer, who can represent a single point of failure. This can be mitigated, e.g., by using threshold blind signatures, where a quorum of authorities must cooperate to create a valid signature (e.g., (Androulaki et al., 2020; Kiayias et al., 2022; Sarencheh et al., 2022)).
@@
Other works reuse the idea of mixing through ring signatures, which conceals the sender’s public key in a group of other public keys—the anonymity set (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Li et al., 2021; Elfadul et al., 2024)). For example, CryptoNote (Van Saberhagen, 2013) employs one-time ring signatures: each transaction uses a fresh key and produces a key image, enabling detection of double spends. The signature hides the actual signer among a group, with anonymity scaling with group size; however, larger rings increase transaction size. Different privacy-preserving crypto-asset systems use ring signatures for obtaining sender anonymity, including Monero [^10] Möser et al. (Möser et al., 2018) showed weaknesses in Monero’s key sampling: $66.09\%$ of transactions do not include any other public key, sampling is not really random, and real inputs can be inferred with $80\%$ accuracy. (Noether et al., 2016), Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024). However, studies show that Monero’s anonymity can be compromised through transaction analysis (e.g., (Möser et al., 2018; Kappos et al., 2018; Kumar et al., 2017)).
@@
To be valid, unspent tokens require proving their ownership. To implement sender anonymity, existing approaches rely on two key tricks. First, the unspent token includes a commitment to its serial number; the serial number is revealed only to mark the token as spent (see Section 4.2.5). Second, the spending operation includes a ZKP that demonstrates knowledge of the opening to a commitment of an unspent coin. Sander and Ta-Shma (Sander and Ta-Shma, 1999) use multiple Merkle trees to track valid Chaumian tokens. To prove validity, the sender provides a ZKP attesting knowledge of a valid coin in the tree and of the coin’s serial number and a hash chain leading to one of Merkle tree roots. Zerocoin (Miers et al., 2013), and its extensions (e.g., (Ben Sasson et al., 2014; Garman et al., 2017)), adopts ZKPs as well: the payer publishes a transaction that includes the recipient’s address, an empty origin address, a ZKP, and the commitment opening. The ZKP prevents linking the sender to any specific commitment; the commitment itself hides the opening using a random value that is never revealed, preventing any association between the two.
Privacy-preserving account-based models mostly rely on ZKPs to obtain sender anonymity (e.g., (Gross et al., 2021; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024)). Specifically, a commitment to the account state is used, with updates including ZKPs to prove correctness. For example, Gross et al. (Gross et al., 2021) use the ideas of Zerocash by committing on the accounts state and by invalidating them on update. The authors suggest using Merkle trees to prove that a transaction proposal refers to an existing commitment in the ledger without pointing to (and thus revealing) it. [^11] For a broader discussion of set accumulators—including Merkle trees as a key example—and their role in enhancing anonymity, see (Loporchio et al., 2023). Platypus (Wüst et al., 2022) uses a model where payer and payee collaborate to exchange with the central bank commitments to their updated accounts, the transaction (amount and serial number), and two ZKPs proving the update’s validity and its link to a prior certified state. If valid the bank updates a public ledger. Conversely, PEReDi (Kiayias et al., 2022) achieves sender anonymity through anonymous channels, an encrypted ledger, and a threshold signature scheme. Sender and receiver submit partial (specular) transactions that update only their respective account states, encrypted using threshold encryption. Solidus (Cecchetti et al., 2017) uses banks as proxies: transactions are not explicitly exchanged between users but among their respective banks.
@@ 4.3.2. Receiver Anonymity
The key techniques for receiver anonymity are one-time destination key or stealth address (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), encrypted address (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)), commitments as output (e.g., (Poelstra, 2016; Wüst et al., 2019)), and updatable public key (e.g., (Fauzi et al., 2019; Papadoulis et al., 2024)). Obtaining receiver anonymity with Chaumian tokens is challenging, as the payee must interact with the bank to deposit the received tokens. Hayes (Hayes, 1990) suggests using one-time aliases to obtain privacy, even though the idea is not detailed for practical usages. A similar idea can be found in CryptoNote (Van Saberhagen, 2013), where the sender generates a one-time destination key by combining a random number with the receiver’s public key—essentially performing a Diffie-Hellman exchange to derive a shared secret. [^12] Because this process requires two elliptic curve keys, a CryptoNote address is twice the size of a standard elliptic curve public key. The sender posts the transaction to a one-time public key with a commitment to a random value. The receiver scans all transactions with his private key, reconstructs the destination key, and accepts the payment if it matches. Sender and receiver never interact directly, but each receiver must scan the entire ledger to find his payments. This approach has been later integrated in other systems, including Monero (Noether et al., 2016), which calls the one-time receiver address as stealth address, Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024) (see Section 3.6).
@@
In Zerocash (Ben Sasson et al., 2014) (unspent token-based ledger), the sender includes an encrypted version of the recipient’s address in the transaction (using asymmetric encryption). Therefore, periodically, each user has to scan the ledger, decrypt transactions using his private key, and identify those intended for him. UTT (Tomescu et al., 2022) (Chaumian token-based system) ensures receiver anonymity using homomorphic commitments and anonymous IBE scheme (Boneh and Franklin, 2003). A coin is a commitment to a tuple containing the owner’s identifier, a serial number issued by the central bank, and its value. To be valid, each coin comes with a (randomizable) signature by the central bank. To make a payment, the sender creates a transaction for the payee, which includes the receiver’s identity commitment, the output value commitment, the encrypted coin details (with the recipient’s key), and ZKPs. The transaction is sent to the bank for validation; if valid, the bank blindly issues new coins for the recipient and records them on a public ledger [^13] The homomorphic properties of the commitment scheme ensures that the bank does not learn identities or amounts, and allows to directly combine the value commitment, identity commitment, and serial number commitment in a single coin commitment.. Recipients periodically scan the ledger to get their coins, to be randomized before adding to the wallet.
@@
The unspent token-based ledger MimbleWimble (Poelstra, 2016) (and PRCash (Wüst et al., 2019)) performs transactions between commitments rather than addresses; this guarantees receiver anonymity and unlinkability.
@@ 4.3.3. Transaction Value Confidentiality
Camenisch et al. (Camenisch et al., 2006) use commitments with verifiable encryption to hide the commitment’s opening. In 2013, Maxwell proposed confidential transactions [^14] https://elementsproject.org/features/confidential-transactions/investigation, which use Pedersen commitments to hide transaction amounts while ensuring correctness (also using range proofs). Monero adopted and extended this technique by combining ring signatures for sender anonymity, obtaining RingCT (Noether et al., 2016). Zerocash (Ben Sasson et al., 2014) uses ZKPs to achieve full transaction privacy: a cryptographic commitment is made to a new coin, including its value, owner, and serial number; then, a ZKP ensures transaction validity without exposing sensitive data. Commitment-based schemes are the most widely adopted to obtain confidential transaction value, especially because their homomorphic versions allow to combine commitments without leaking sensitive data. To prevent users from creating negative or infinite amounts, range proofs prove that a committed value is greater than zero and within a valid range.
@@
In PEReDi (Kiayias et al., 2022), senders and receivers update their accounts by submitting encrypted transactions to several maintainers who operate a distributed ledger. Threshold ElGamal encryption (Elgamal, 1985) requires a quorum of maintainer to cooperate to decrypt transaction details. Authorities can trace both transactions and identities. PARScoin (Sarencheh et al., 2025) extends this approach to introduce an offline payment protocol that does not require interaction between sender and receiver. In Solidus (Cecchetti et al., 2017), Publicly Verifiable Oblivious RAMs (PVORMs) enable banks to update the accounts of their clients without revealing exactly which accounts are being updated. PVORM provides a ZKP demonstrating that the updates are correct with respect to the transaction triggering them.
A few solutions implement privacy-preserving digital payments using smart contracts on public blockchains. To ensure confidentiality, transactions are encrypted with the sender’s public key and supplemented with a NIZK proof for public verifiability (e.g., (Ma et al., 2021; Bünz et al., 2020)). For example, Zether (Bünz et al., 2020) uses smart contracts to define a privacy-preserving tokens that can be exchanged between public keys associated with a confidential balance. They use ElGamal encryption, which has homomorphic properties, for hiding each account’s balance. Zether introduces the $\Sigma$-Bullets (an enhancement of the Bulletproofs range ZKP) to efficiently prove statements over the encrypted transfer value and the new sender balance. However, Zether currently allows only one anonymous transfer per epoch (to prevent double spending), and execution costs (namely, gas costs on Ethereum) are very high (Almashaqbeh and Solomon, 2022). Similarly, DSC (Ma et al., 2021) optimizes privacy within Ethereum-style smart contract environments, but does not support anonymity: it uses ZKPs and homomorphic encryption to hide balances and transferred amounts. User balances are stored on the ledger in encrypted form.
@@ 4.3.4. Network Anonymity
Network anonymity concerns the protection of sender and receiver IP addresses, which could otherwise reveal real-world identities. Most of the existing approaches do not investigate (or even mention) the network anonymity feature. Few works mention it but consider this issue as out of scope of their investigation (e.g., (Cecchetti et al., 2017; Baudet et al., 2023; Tomescu et al., 2022; Wüst et al., 2022)). Monero (Noether et al., 2016) and Traceable Monero (Li et al., 2021) explicitly support network anonymity by using TOR (The Onion Router [^15] https://www.torproject.org/) and I2P (Invisible Internet Project [^16] https://geti2p.net/en/) to hide the real IP address (Shi et al., 2024). While TOR routes traffic through multiple relays to obscure the user’s IP address, I2P is a network optimized for anonymous peer-to-peer communication, using fully decentralized routing and encryption to enable different anonymous services. Moreover, Monero uses Dandelion++ (Fanti et al., 2018), an improvement of Dandelion (Bojja Venkatakrishnan et al., 2017), which prevents malicious attackers from linking transactions with their source IPs. Dandelion++ uses randomization mechanisms for message forwarding, which happens in two phases: stem (anonymization) and fluff (broadcast). In the stem phase, a transaction is first relayed privately across a small, randomly chosen subset of nodes; each node probabilistically decides whether to continue forwarding the transaction along the stem or to switch to the second phase. In the fluff phase, the node broadcasts to the entire network using traditional flood propagation protocol.
4.3.5. Linkability and Unlinkability
Unlinkability ensures that transactions cannot be correlated or traced back to the same user (other definitions in (Chaum, 1983; Amarasinghe et al., 2019; Zhang, 2023)). Unlinkability can refer to the sender, receiver, or both. A few works break this property into two facets, internal and external. For instance, CoinJoin (Maxwell, 2013) does not preserve internal unlinkability as each participant in the group that mixes transactions can link entities within the group.
Obtaining unlinkability in public ledgers is challenging as linking between addresses in crypto-asset systems stems from different patterns (e.g., (Ron and Shamir, 2013; Chan and Olmsted, 2017; Béres et al., 2021; Victor, 2020; Zhong and Mueen, 2024; Lu et al., 2022; Ermilov et al., 2017; Kalodner et al., 2020)). A non exhaustive list of patterns and heuristics to link addresses is the following.

### v3

@@ 4.3. Privacy Features
Privacy in digital payment systems is often defined through several sub-properties, including participant anonymity, value confidentiality, plausible deniability, and unlinkability (e.g., (Almashaqbeh and Solomon, 2022; Amarasinghe et al., 2019; Bernal Bernabe et al., 2019; Genkin et al., 2018)). The privacy mechanisms discussed in this section are largely shared across generations; differences arise primarily in how these mechanisms are combined and integrated with system-level objectives, with techniques such as encryption, ZKPs, and credentials [^17] By credentials we refer to cryptographic constructions—typically based on digital signatures combined with ZKPs—that bind attributes or rights to a user and enable selective disclosure without revealing unnecessary information. enabling selective disclosure and auditability.
@@ 4.3.1. Sender Anonymity
Asymmetric anonymity indicates that only one of the participants’ identities (sender or receiver) is protected (e.g., (Chaum, 1983; Sander and Ta-Shma, 1999; Miers et al., 2013; Dold, 2019; Zhang et al., 2019; BIS Innovation Hub, 2023; Goodell et al., 2023; Yu et al., 2024; Androulaki and Karame, 2014)). Sender anonymity focuses only on hiding the payer’s identity, usually using the following techniques: blind signatures (e.g., (Chaum, 1983; Dold, 2019; Goodell et al., 2023; Liu et al., 2022)) and their variants—one-time (e.g., (Hayes, 1990)), threshold (e.g., (Sarencheh et al., 2025; Kiayias et al., 2022; Androulaki et al., 2020)), and randomizable (e.g., (Tomescu et al., 2022; Wüst et al., 2019))—, ring signatures (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), group signatures (e.g., (Zhang et al., 2019)), and commitments with ZKPs (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)).
@@
Chaum (Chaum, 1983) proposes blind signatures to unlink the sender request of token issuance and its later spending. To reveal the sender on double spending, Hayes (Hayes, 1990) introduces one-time blind signatures. However, since each transaction requires a new signing key, storage and computation management become cumbersome; moreover, one-time signatures often require large signature size and expensive verification, making them impractical for high-throughput systems. UTT (Tomescu et al., 2022) implements blind signatures using randomizable signatures (Pointcheval and Sanders, 2016), which allow signatures to be re-randomized using fresh randomness without compromising their validity. In particular, each coin consists of a commitment signed by the central bank, where the commitment binds the payer’s public identifier, the coin value, and a random nonce (to ensure hiding). All these schemes rely on a central issuer, which represents a potential single point of failure. This can be mitigated, e.g., by using threshold blind signatures, where a quorum of authorities must cooperate to create a valid signature (e.g., (Androulaki et al., 2020; Kiayias et al., 2022; Sarencheh et al., 2022)).
@@
Other works reuse the idea of mixing through ring signatures, which conceals the sender’s public key within a group of other public keys—the anonymity set (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Li et al., 2021; Elfadul et al., 2024)). For example, CryptoNote (Van Saberhagen, 2013) employs one-time ring signatures: each transaction uses a fresh key and produces a key image, enabling detection of double spends. Different privacy-preserving crypto-asset systems use ring signatures for obtaining sender anonymity, including Monero [^18] Möser et al. (Möser et al., 2018) showed weaknesses in Monero’s key sampling: $66.09\%$ of transactions do not include any other public key, sampling is not really random, and real inputs can be inferred with $80\%$ accuracy. (Noether et al., 2016), Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024). However, studies show that Monero’s anonymity can be compromised through transaction analysis (e.g., (Möser et al., 2018; Kappos et al., 2018; Kumar et al., 2017)).
@@
To be valid, unspent tokens require proving their ownership. To implement sender anonymity, existing approaches rely on two key techniques. First, the unspent token includes a commitment to its serial number; the serial number is revealed only to mark the token as spent (see Section 4.2.5). Second, the spending operation includes a ZKP that demonstrates knowledge of the opening to a commitment of an unspent coin. Sander and Ta-Shma (Sander and Ta-Shma, 1999) use multiple Merkle trees to track valid Chaumian tokens. To prove validity, the sender provides a ZKP attesting knowledge of a valid coin in the tree and of the coin’s serial number and a hash chain leading to one of Merkle tree roots. Zerocoin (Miers et al., 2013), and its extensions (e.g., (Ben Sasson et al., 2014; Garman et al., 2017)), adopts ZKPs as well: the payer publishes a transaction that includes the recipient’s address, an empty origin address, a ZKP, and the commitment opening. The ZKP prevents linking the sender to any specific commitment; the commitment itself hides the opening using a random value that is never revealed, preventing any association between the two.
Privacy-preserving account-based models mostly rely on ZKPs to obtain sender anonymity (e.g., (Gross et al., 2021; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024)). Specifically, a commitment to the account state is used, with updates including ZKPs to prove correctness. For example, Gross et al. (Gross et al., 2021) use the ideas of Zerocash by committing on the accounts state and by invalidating them on update. The authors suggest using Merkle trees to prove that a transaction proposal refers to an existing commitment in the ledger without pointing to (and thus revealing) it. [^19] For a broader discussion of set accumulators—including Merkle trees as a key example—and their role in enhancing anonymity, see (Loporchio et al., 2023). Platypus (Wüst et al., 2022) uses a model where payer and payee collaborate to exchange with the central bank commitments to their updated accounts, the transaction (amount and serial number), and two ZKPs proving the update’s validity and its link to a prior certified state. If valid the bank updates a public ledger. Conversely, PEReDi (Kiayias et al., 2022) achieves sender anonymity through anonymous channels, an encrypted ledger, and a threshold signature scheme. Sender and receiver submit partial (specular) transactions that update only their respective account states, encrypted using threshold encryption. Solidus (Cecchetti et al., 2017) is a hierarchical system where banks act as proxies, managing on-chain transactions on behalf of users and ensuring transaction-graph confidentiality. [^20] While banks can identify their respective clients, other entities only learn the identities of the banks.
@@ 4.3.2. Receiver Anonymity
The key techniques for receiver anonymity are one-time destination keys or stealth address (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024; Zhang et al., 2024)), encrypted address (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)), commitments as output (e.g., (Poelstra, 2016; Wüst et al., 2019)), and updatable public key (e.g., (Fauzi et al., 2019; Papadoulis et al., 2024)). Obtaining receiver anonymity with Chaumian tokens is challenging, as the payee must interact with the bank to deposit the received tokens. Hayes (Hayes, 1990) suggests using one-time aliases to obtain privacy, even though the idea is not detailed for practical usages. A similar idea can be found in CryptoNote (Van Saberhagen, 2013), where the sender generates a one-time destination key by combining a random number with the receiver’s public key—essentially performing a Diffie-Hellman exchange to derive a shared secret. [^21] Because this process requires two elliptic curve keys, a CryptoNote address is twice the size of a standard elliptic curve public key. The sender posts the transaction to a one-time public key with a commitment to a random value. The receiver scans all transactions with his private key, reconstructs the destination key, and accepts the payment if it matches. Sender and receiver never interact directly, but each receiver must scan the entire ledger to find his payments. This approach has been later integrated in other systems, including Monero (Noether et al., 2016), which calls the one-time receiver address as stealth address, Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024) (see Section 3.6).
@@
In Zerocash (Ben Sasson et al., 2014) (unspent token-based ledger), the sender includes an encrypted version of the recipient’s address in the transaction (using asymmetric encryption). Therefore, periodically, each user has to scan the ledger, decrypt transactions using his private key, and identify those intended for him. This prevents external observers from linking payments to specific recipients while allowing recipients to identify incoming transactions without revealing their addresses publicly. UTT (Tomescu et al., 2022) (Chaumian token-based system) ensures receiver anonymity using homomorphic commitments and anonymous IBE scheme (Boneh and Franklin, 2003). A coin is a commitment to a tuple containing the owner’s identifier, a serial number issued by the central bank, and its value. To be valid, each coin comes with a (randomizable) signature by the central bank. To make a payment, the sender creates a transaction for the payee, which includes the receiver’s identity commitment, the output value commitment, the encrypted coin details (with the recipient’s key), and ZKPs. The transaction is sent to the bank for validation; if valid, the bank blindly issues new coins for the recipient and records them on a public ledger [^22] The homomorphic properties of the commitment scheme ensures that the bank does not learn identities or amounts, and allows to directly combine the value commitment, identity commitment, and serial number commitment in a single coin commitment.. Recipients periodically scan the ledger to get their coins, to be randomized before adding to the wallet.
@@
The unspent token-based ledger MimbleWimble (Poelstra, 2016) (and PRCash (Wüst et al., 2019)) performs transactions between commitments rather than addresses; this guarantees both receiver anonymity and unlinkability.
@@ 4.3.3. Transaction Value Confidentiality
Camenisch et al. (Camenisch et al., 2006) use commitments with verifiable encryption to hide the commitment’s opening. In 2013, Maxwell proposed Confidential Transactions [^23] https://elementsproject.org/features/confidential-transactions/investigation, which use Pedersen commitments to hide transaction amounts while ensuring correctness via range proofs (e.g., Bulletproofs). Monero combines this technique with ring signatures obtaining RingCT (Noether et al., 2016). Zerocash (Ben Sasson et al., 2014) uses ZKPs to achieve full transaction privacy: a cryptographic commitment is made to a new coin, including its value, owner, and serial number; then, a ZKP ensures transaction validity without exposing sensitive data. Commitment-based schemes are the most widely adopted to obtain confidential transaction value, especially because their homomorphic versions allow to combine commitments without leaking sensitive data. To prevent users from creating negative or infinite amounts, range proofs prove that a committed value is greater than zero and within a valid range.
@@
In PEReDi (Kiayias et al., 2022), senders and receivers update their accounts by submitting encrypted transactions to several maintainers who operate a distributed ledger. Threshold ElGamal encryption (Elgamal, 1985) requires a quorum of maintainers to cooperate to decrypt transaction details. Authorities can trace both transactions and identities. PARScoin (Sarencheh et al., 2025) extends this approach to introduce an offline payment protocol that does not require interaction between sender and receiver. In Solidus (Cecchetti et al., 2017), Publicly Verifiable Oblivious RAMs (PVORMs) enable banks to update the accounts of their clients without revealing exactly which accounts are being updated. PVORM provides a ZKP demonstrating that the updates are correct with respect to the transaction triggering them.
A few solutions implement privacy-preserving digital payments using smart contracts on public blockchains. To ensure confidentiality, transactions are encrypted with the sender’s public key and supplemented with a NIZK proof for public verifiability (e.g., (Ma et al., 2021; Bünz et al., 2020)). For example, Zether (Bünz et al., 2020) uses smart contracts to define privacy-preserving tokens that can be exchanged between public keys associated with a confidential balance. They use ElGamal encryption, which has homomorphic properties, to hide each account’s balance. Zether introduces the $\Sigma$-Bullets (an enhancement of the Bulletproofs range ZKP) to efficiently prove statements over the encrypted transfer value and the new sender balance. However, Zether currently allows only one anonymous transfer per epoch (to prevent double spending), and execution costs (namely, gas costs on Ethereum) are very high (Almashaqbeh and Solomon, 2022). Similarly, DSC (Ma et al., 2021) optimizes privacy within Ethereum-style smart contract environments, but does not support anonymity: it uses ZKPs and homomorphic encryption to hide balances and transferred amounts. User balances are stored on the ledger in encrypted form. The homomorphic properties allows validators to update the balance in ciphertext without the need of decryption.
@@ 4.3.4. Network Anonymity
Network anonymity concerns the protection of sender and receiver IP addresses, which could otherwise reveal real-world identities. Most of the existing approaches do not investigate (or even mention) the network anonymity feature. Few works mention it but consider this issue as out of scope of their investigation (e.g., (Cecchetti et al., 2017; Baudet et al., 2023; Tomescu et al., 2022; Wüst et al., 2022)). They consider this feature as orthogonal hence that can be achieved by implementing the digital payment system on top of a private overlay network. Monero (Noether et al., 2016) and Traceable Monero (Li et al., 2021) explicitly support network anonymity by using TOR (The Onion Router [^24] https://www.torproject.org/) and I2P (Invisible Internet Project [^25] https://geti2p.net/en/) to hide the real IP address (Shi et al., 2024). While TOR routes traffic through multiple relays to obscure the user’s IP address, I2P is a network optimized for anonymous peer-to-peer communication, using fully decentralized routing and encryption to enable different anonymous services. Moreover, Monero uses Dandelion++ (Fanti et al., 2018), an improvement of Dandelion (Bojja Venkatakrishnan et al., 2017), which prevents attackers from linking transactions to their source IP addresses. Dandelion++ uses randomization mechanisms for message forwarding, which happens in two phases: stem (anonymization) and fluff (broadcast). In the stem phase, a transaction is first relayed privately across a small, randomly chosen subset of nodes; each node probabilistically decides whether to continue forwarding the transaction along the stem or to switch to the second phase. In the fluff phase, the node broadcasts to the entire network using traditional flood propagation protocol.
4.3.5. Unlinkability
Unlinkability ensures that transactions cannot be correlated or traced back to the same user (other definitions in (Chaum, 1983; Amarasinghe et al., 2019; Zhang, 2023)). It can refer to the sender, receiver, or both. Obtaining unlinkability in public ledgers is challenging as links between addresses in crypto-asset systems stems from different patterns (e.g., (Ron and Shamir, 2013; Chan and Olmsted, 2017; Béres et al., 2021; Victor, 2020; Zhong and Mueen, 2024; Lu et al., 2022; Ermilov et al., 2017; Kalodner et al., 2020)); a non-exhaustive list follows.